# Artin–Schreier quandles of involutions in absolute Galois groups

### 12 Mar 2024 v1

@@ 2 Involutions and quandles
Let us set $\mathrm{FQ}^{\wedge}_{2}(S)=\operatorname{Inv}(\mathrm{F}^{\wedge}_{2}(S))$. This is a pro-finite quandle: the involutions are a closed subset of the topological space $\mathrm{F}^{\wedge}_{2}(S)$ as they are the image of the continuous map $\mathrm{F}^{\wedge}_{2}(S)\times S\to\mathrm{F}^{\wedge}_{2}(S)$ that sends $(g,s)$ to $gsg^{-1}$. The involutory quandle $\mathrm{FQ}^{\wedge}_{2}(S)$ comes with a continuous map from $S$ because the image of the canonical map $S\to\mathrm{F}^{\wedge}_{2}(S)$ lies inside $\mathrm{FQ}^{\wedge}_{2}(S)$.
Clearly, any continuous morphism $\mathrm{FQ}^{\wedge}_{2}(S)\to Q$ from $\mathrm{FQ}^{\wedge}_{2}(S)$ into a pro-finite involutory quandle $Q$ restricts to a continuous map $S\to Q$. Conversely, if $\varphi\colon S\to Q$ is any such map, we can extend it functorially to a continuous morphism $\mathrm{F}^{\wedge}_{2}(\varphi)\colon\mathrm{F}^{\wedge}_{2}(S)\to\mathrm{F}^{\wedge}_{2}(Q)$ of groups. If we now have an element in $\mathrm{FQ}^{\wedge}_{2}(S)=\operatorname{Inv}(\mathrm{F}^{\wedge}_{2}(S))$, it will be conjugate to an element in $S$, so that we can write it as $gsg^{-1}$ with $g$ in $\mathrm{F}^{\wedge}_{2}(S)$ and $s$ in $S$. In fact, as the centraliser of $s$ in $\mathrm{F}^{\wedge}_{2}(S)$ is $\{e,s\}$, any other such expression $hsh^{-1}=gsg^{-1}$ will have $h\in\{g,gs\}$. Therefore, if we try to define our extension of $\varphi$ to a quandle morphism $\mathrm{FQ}^{\wedge}_{2}(S)\to Q$ by sending $gsg^{-1}$ to $\mathrm{F}^{\wedge}_{2}(\varphi)(g)\varphi(s)$, where $\mathrm{F}^{\wedge}_{2}(Q)$ acts on $Q$ via left-multiplications, we only have to check that $\mathrm{F}^{\wedge}_{2}(\varphi)(gs)$ acts exactly the same as $\mathrm{F}^{\wedge}_{2}(\varphi)(gs)$ does on $\varphi(s)$. But that is easy: as $\mathrm{F}^{\wedge}_{2}(\varphi)$ is a morphism of groups, we have $\mathrm{F}^{\wedge}_{2}(\varphi)(gs)=\mathrm{F}^{\wedge}_{2}(\varphi)(g)\mathrm{F}^{\wedge}_{2}(\varphi)(s)$, and $\mathrm{F}^{\wedge}_{2}(\varphi)(s)=\varphi(s)$ acts trivially on itself by the quandle axioms. The preceding two constructions are inverse to each other. ∎
@@ 3 Artin–Schreier quandles
We recall that an ordering of a field $F$ is given by a subset $P\subset F$ of elements that is closed under addition and multiplication and satisfies $P\cap-P=\{0\}$ and $P\cup-P=F$, where $-P=\{a\in F\,|\,-a\in P\}$. It follows that all sums of squares are in $P$. A field is formally real if and only if we can order it. A Euclidean field has a unique ordering, and in that ordering, the set $P$ is the set of squares. A field is real closed if, in addition, each polynomial of odd degree with coefficients in $F$ has at least one root in $F$.
The real spectrum of a field $F$ is the set $\operatorname{Ord}(F)$ of orderings of $F$ equipped with the Harrison topology: given any element $a\not=0$ in $F$, the set of orderings for which $a$ is positive is a generating open set. With this topology, the real spectrum is a pro-finite space.
@@
Given any ordering $P$ of $F$, the associated morphism sends a quadratic form to its $P$–signature—the difference of the numbers of positive and negative diagonal entries, where ‘positive’ and ‘negative’ is determined by the ordering $P$. Another way to think about this: an ordering determines an embedding into a real closed field so that we can use functoriality and the fact that the Witt ring of a real closed field is canonically isomorphic to $\mathbb{Z}$ through the signature. ∎
@@
The kernel $I$ of the parity homomorphism $\mathrm{W}(F)\to\mathbb{Z}/2$ defines the $I$–adic filtration on the Witt ring $\mathrm{W}(F)$. The associated graded ring is $\operatorname{gr}\mathrm{W}(F)=\mathrm{W}/I\oplus I/I^{2}\oplus I^{2}/I^{3}\oplus\dots$. It is a fact that $\operatorname{gr}\mathrm{W}(F)$ is a graded commutative algebra over $\mathbb{F}_{2}$, and it is isomorphic to mod $2$ Milnor K-theory and mod $2$ Galois cohomology via Milnor’s conjecture, which is now a theorem [43]. Therefore, we can recover $\operatorname{gr}\mathrm{W}(F)$ from the absolute Galois group of $F$. Can we recover $\mathrm{W}(F)$ itself from a refinement of the absolute Galois group? This was a non-trivial question because there are examples of fields with isomorphic $\operatorname{gr}\mathrm{W}(F)$’s but non-isomorphic $\mathrm{W}(F)$’s (see [44]); it has been answered in [39, 40].
@@
If $E|F$ is a Galois extension with Galois group $G$, there is a map $\mathrm{A}(G)\to\mathrm{W}(F)$ from the Burnside ring of the Galois group to the Witt ring of the ground field that sends $G/H$ to $(E^{H},\operatorname{tr})$, where $\operatorname{tr}$ is the usual trace form $(a,b)\mapsto\operatorname{tr}_{E|F}(ab)$. It is surjective if $E=F(\sqrt{a}\,|\,a\in F)$.
@@
Given a conjugacy class $[\sigma]$ of an involution $\sigma$ in $\operatorname{Gal}(F)$, we have the ordering $P[\sigma]=\{a\in F\,|\,\sigma(\sqrt{a})=\sqrt{a}\}$. Note that the condition $\sigma(\sqrt{a})=\sqrt{a}$ does not depend on the choice of the square root $\sqrt{a}$ of $a$. ∎
@@
Let $F$ be a field and $\operatorname{Gal}(F)$ be its absolute Galois group. From the work of Artin and Schreier cited in the introduction, it is known that the group $\operatorname{Gal}(F)$, even though pro-finite, can only have non-trivial elements of finite order if $\operatorname{char}(F)=0$, and all of these elements necessarily have order $2$, i.e., they are involutions. We refer to [28] for an interesting discussion of torsion in absolute Galois groups from a different angle. The following lemma collects useful group theoretic statements about involutions in absolute Galois groups.
@@
If $\tau$ is a field automorphism of $\operatorname{Fix}(\sigma)$ that fixes $F$, and $a\in\operatorname{Fix}(\sigma)$ with $\tau(a)\not=a$, we have $a<\tau(a)$, say. Then $a<\tau(a)<\tau^{2}(a)<\tau^{3}(a)<\dots$, contradicting the algebraicity of $a$: the $\tau$–orbit of $a$ must be finite. The argument for $a>\tau(a)$ is similar, and we have $\tau(a)=a$ for all $a\in\operatorname{Fix}(\sigma)$, and this shows $\operatorname{Aut}(\operatorname{Fix}(\sigma)|F)=\{\operatorname{id}\}$.
If an element $\tau\in\operatorname{Gal}(F)$ lies in the centraliser of $\sigma$, then $\tau$ sends $\operatorname{Fix}(\sigma)$ into itself. From what we have shown before, it follows that $\tau$ is the identity on $\operatorname{Fix}(\sigma)$ and lies in $\operatorname{Gal}(\operatorname{Fix}(\sigma))=\langle\sigma\rangle$. The last statement is simply another reformulation of the preceding statement. ∎
Later on, we shall need the following stronger statement about involutions in absolute Galois groups.
@@
By Artin–Schreier, we already know that every element of the group $G$ must have order $1$ or $2$. It follows from group theory that $G$ is an elementary abelian $2$–group, that is $G\cong(\mathbb{Z}/2)^{n}$ for some $n$. We show $n=1$ by showing that $\operatorname{Gal}(F)$ cannot contain a subgroup of the form $\mathbb{Z}/2\times\mathbb{Z}/2$: Given an involution $\sigma\not=\operatorname{id}$, Lemma 3.5 says that its centraliser within $\operatorname{Gal}(F)$ is the subgroup $\langle\sigma\rangle$ generated by $\sigma$. Therefore, there can be no other element in the group $\operatorname{Gal}(F)$ that commutes with $\sigma$. ∎
@@
For any field $F$, the set $\operatorname{Inv}(\operatorname{Gal}(F))$ is a closed subspace of the absolute Galois group $\operatorname{Gal}(F)$. Therefore, it is pro-finite as well.
@@
Let $F$ be a field. The Artin–Schreier quandle $\mathrm{AS}(F)$ of $F$ is the pro-finite involutory quandle
@@
The Artin–Schreier quandle $\mathrm{AS}(F)$ of a field $F$, just as its absolute Galois group $\operatorname{Gal}(F)$, depends on the choice of an algebraic closure of $F$. If $\omega\colon\Omega_{1}\to\Omega_{2}$ is an $F$–isomorphism of algebraic closures of $F$, then $\sigma\mapsto\omega\sigma\omega^{-1}$ is an isomorphism $\operatorname{Aut}(\Omega_{1}|F)\to\operatorname{Aut}(\Omega_{2}|F)$ and this isomorphism restricts to an isomorphism between the Artin–Schreier quandles computed using $\Omega_{1}$ and $\Omega_{2}$, respectively. As usual, we will suppress the choice of an algebraic closure from the notation.
@@
Before we come to examples, let us note the following algebraic result about a special property of Artin–Schreier quandles.
@@
The ‘if’ direction is part of the quandle axioms, which are clear by now. For the ‘only if’ direction, note that $\sigma\rhd\tau=\tau$ is equivalent to $\sigma\tau=\tau\sigma$, so that $\tau$ must be in the centraliser of $\sigma$. It cannot be the identity, by definition of $\mathrm{AS}(F)$, so the result follows from Lemma 3.6. ∎
@@
A field $F$ is known to be formally real if and only if its absolute Galois group $\operatorname{Gal}(F)$ contains an involution. ∎
@@
Proposition 3.12 implies that it is not interesting—for our purposes—to consider fields that are not formally real. Therefore, unless otherwise stated, we can assume that we are dealing with formally real fields.
@@
A field $F$ is known to be real closed if and only if its absolute Galois group has order $2$. ∎
@@
The field $\mathbb{R}$ of real numbers is real closed, and the unique element in the involutory quandle $\mathrm{AS}(\mathbb{R})=\{\gamma\}$ is complex conjugation $\gamma\colon z\mapsto\overline{z}$ inside the absolute Galois group $\operatorname{Gal}(\mathbb{R})=\operatorname{Aut}(\mathbb{C}|\mathbb{R})=\{\operatorname{id},\gamma\}$ of order $2$. More generally, let $F$ be a number field that is Galois over $\mathbb{Q}$ and not totally real. We can embed $F$ into $\mathbb{C}$, and apply complex conjugation. This gives an involution in $\operatorname{Gal}(F)$, but this involution may depend on the embedding—unless $F$ is a totally imaginary quadratic extension of a totally real number field (i.e., a CM field), such as $\mathbb{Q}(\sqrt{-1})$.
@@
A bijection exists between the set $\mathrm{AS}(F)$ and the set of real closed subfields of the algebraic closure used to define the absolute Galois group. An involution $\sigma$ corresponds to its fixed field $\operatorname{Fix}(\sigma)$.
@@
If the Artin–Schreier quandle $\mathrm{AS}(F)$ contains more than one element, then the absolute Galois group $\operatorname{Gal}(F)$ contains two different involutions, say $\sigma\not=\tau$. Consider the morphism $\mathrm{F}_{2}\{s,t\}\to\operatorname{Gal}(F)$ of groups that sends the generator $s$ to $\sigma$ and $t$ to $\tau$. Suppose there were a non-trivial kernel. This kernel cannot contain $s$ or $t$ by definition, and it cannot contain any other involution, as these are each conjugate to either $s$ or $t$ inside $\mathrm{F}_{2}\{s,t\}$. It follows that the kernel is a subgroup of the infinite cyclic subgroup generated by $st$, so that the image is a finite (dihedral) subgroup of $\operatorname{Gal}(F)$, and it contains two different involutions $\sigma\not=\tau$. But the finite subgroups of $\operatorname{Gal}(F)$ have order at most $2$ by Artin–Schreier—a contradiction. Hence, the morphism $\mathrm{F}_{2}\{s,t\}\to\operatorname{Gal}(F)$ is injective, and $\operatorname{Gal}(F)$ contains infinitely many involutions because this already holds for the infinite dihedral group $\mathrm{F}_{2}\{s,t\}$, and these lie in $\mathrm{AS}(F)$. ∎

### 08 Apr 2026 v2

@@ 2 Involutions and quandles
This is a profinite quandle: the involutions form a closed subset of the topological space $\mathrm{F}^{\wedge}_{2}(S)$ as they are the image of the continuous map $\mathrm{F}^{\wedge}_{2}(S)\times S\to\mathrm{F}^{\wedge}_{2}(S)$ that sends $(g,s)$ to $gsg^{-1}$. The involutory quandle $\mathrm{FQ}^{\wedge}_{2}(S)$ comes with a natural continuous map from $S$ because the image of the canonical map $S\to\mathrm{F}^{\wedge}_{2}(S)$ lies inside $\mathrm{FQ}^{\wedge}_{2}(S)$. So, any continuous morphism $\mathrm{FQ}^{\wedge}_{2}(S)\to Q$ from $\mathrm{FQ}^{\wedge}_{2}(S)$ into a profinite involutory quandle $Q$ restricts to a continuous map $S\to Q$.
Conversely, if $\varphi\colon S\to Q$ is any continuous map, we can extend it functorially to a continuous morphism $\mathrm{F}^{\wedge}_{2}(\varphi)\colon\mathrm{F}^{\wedge}_{2}(S)\to\mathrm{F}^{\wedge}_{2}(Q)$ of groups, and this restricts to a continuous morphism $\mathrm{FQ}^{\wedge}_{2}(\varphi)\colon\mathrm{FQ}^{\wedge}_{2}(S)\to\mathrm{FQ}^{\wedge}_{2}(Q)$ of involutory quandles. We get a continuous morphism $\mathrm{FQ}^{\wedge}_{2}(S)\to Q$ by composition of this morphism $\mathrm{FQ}^{\wedge}_{2}(\varphi)$ with the continuous orbit map $\mathrm{FQ}^{\wedge}_{2}(Q)\to Q$, where $Q$ is identified with the space of orbits in the subspace $\mathrm{FQ}^{\wedge}_{2}(Q)=\operatorname{Inv}(\mathrm{F}^{\wedge}_{2}(Q))$ for the continuous action of the group $\mathrm{F}^{\wedge}_{2}(Q)$ on itself via conjugation. The preceding two constructions are inverse to each other.
The final statement is immediate from what we have already used in the previous paragraph of the proof: The composition $S\longrightarrow\mathrm{FQ}^{\wedge}_{2}(S)\longrightarrow\mathrm{FQ}^{\wedge}_{2}(S)/\mathrm{F}^{\wedge}_{2}(S)$ is a homeomorphism, and we can use its inverse to obtain the desired map. ∎
Remark 2.17.
We remark that any free quandle generates a free group: Ignoring topologies for a moment, the enveloping group $\mathrm{Gr}(Q)$ of a quandle $Q$ is, by definition, the value of the left-adjoint $\mathrm{Gr}$ to the forgetful functor from the category of groups to the category of quandles. To construct it, up to isomorphism, we can start with the free group $\mathrm{F}(Q)$ on the set $Q$, with $f(q)$ the basis element corresponding to an element $q\in Q$, and pass to the quotient by the relations
$f(x\rhd y)=f(x)f(y)f(x)^{-1}.$ (2.6)
As a left adjoint, the functor $\mathrm{Gr}$ preserves colimits, and this shows that we have an isomorphism $\mathrm{Gr}(\operatorname{FQ}(S))\cong\mathrm{F}(S)$. The construction can be adapted to profinite quandles $Q$. Ignoring the quandle structure on $Q$ for a moment, the profinite space $Q$ freely generates the profinite group $\mathrm{F}^{\wedge}(Q)$, and $\mathrm{Gr}^{\wedge}(Q)$ is the quotient of it by the closed normal subgroup generated by the relations (2.6). Similarly, the enveloping group of the quandle $\mathrm{FQ}^{\wedge}_{2}(S)$ is canonically isomorphic to the group $\mathrm{F}^{\wedge}_{2}(S)$ for formal reasons; both sides satisfy the same universal property. All that is to point out that, in contrast, the proof of Theorem 4.1 will go the other way. We will argue for the freeness of a quandle from the freeness of its enveloping group, and such an implication is not just a categorical formality. Any argument to deduce the freeness of a quandle $Q$ from the freeness of its enveloping group has to be non-formal.
Remark 2.18.
We shall also use the fact that the map $\mathrm{FQ}^{\wedge}_{2}(S)\to\mathrm{F}^{\wedge}_{2}(S)$ is injective. This is not automatic, either. There are quandles $Q$ for which the canonical map from $Q$ into the enveloping group is not injective. For example, if $Q=\{x,y,z\}$ is the involutory quandle with $\lambda_{x}$ transposing $y$ and $z$, and $\lambda_{y}=\lambda_{z}=\operatorname{id}_{Q}$, then $y\rhd x=x$ implies that $x$ commutes with $y$ in the enveloping group, and $x\rhd y=z$ implies that $x$ conjugates $y$ into $z$ in the enveloping group. Therefore, the elements $y$ and $z$ must become equal in the enveloping group.
@@ 3 Artin–Schreier quandles
We recall that an ordering of a field $F$ is given by a subset $P\subset F$ of elements that is closed under addition and multiplication and satisfies $P\cap-P=\{0\}$ and $P\cup-P=F$, where $P=\{a\in F\mid-a\in P\}$. It follows that all sums of squares are in $P$. A field is formally real if and only if we can order it. A Euclidean field is a field for which the set of squares is an ordering. In a Euclidean field, the ordering is necessarily unique. A field is real closed if, in addition, each polynomial of odd degree with coefficients in $F$ has at least one root in $F$.
The real spectrum of a field $F$ is the set $\operatorname{Ord}(F)$ of orderings of $F$ equipped with the Harrison topology: given any element $a\not=0$ in $F$, the set of orderings for which $a$ is positive is a generating open set. With this topology, the real spectrum is a profinite space.
@@
Given any ordering $P$ of $F$, the associated morphism sends a quadratic form to its PP–signature—the difference of the numbers of positive and negative diagonal entries, where the meaning of ‘positive’ and ‘negative’ is determined by the ordering $P$. Another way to think about this: an ordering determines an embedding into a real closed field so that we can use functoriality and the fact that the Witt ring of a real closed field is canonically isomorphic to $\mathbb{Z}$ through the signature. We refer to [33, Thm. I.2.3] for more detail. ∎
@@
The kernel $I$ of the parity homomorphism $\mathrm{W}(F)\to\mathbb{Z}/2$ defines the II–adic filtration on the Witt ring $\mathrm{W}(F)$, and the associated graded ring is then $\operatorname{gr}\mathrm{W}(F)=\mathrm{W}(F)/I\oplus I/I^{2}\oplus I^{2}/I^{3}\oplus\dots$. It is a fact that $\operatorname{gr}\mathrm{W}(F)$ is a graded commutative algebra over $\mathbb{F}_{2}$, and it is isomorphic to mod $2$ Milnor K-theory and mod $2$ Galois cohomology via Milnor’s conjecture, which is now a theorem [48]. Therefore, we can recover $\operatorname{gr}\mathrm{W}(F)$ from the absolute Galois group of $F$. Can we recover $\mathrm{W}(F)$ itself from a refinement of the absolute Galois group? This was a non-trivial question because there are examples of fields with isomorphic $\operatorname{gr}\mathrm{W}(F)$’s but non-isomorphic $\mathrm{W}(F)$’s (see [49]); it has been answered in the affirmative by Mináč and Spira in [43, 44]. The authors show that the Witt ring determines and is determined (modulo a small restriction) by a specific quotient of the absolute Galois group (see [44, Thm. 3.8] in particular).
@@
Half a century ago, Dress [15] found a relation between Witt rings and Burnside rings: If $E|F$ is a Galois extension with Galois group $G$, then there is a map $\mathrm{A}(G)\to\mathrm{W}(F)$ from the Burnside ring $\mathrm{A}(G)$ of the Galois group to the Witt ring of the ground field that sends $G/H$ to $(E^{H},\operatorname{tr})$, where $\operatorname{tr}$ is the usual trace form $(a,b)\mapsto\operatorname{tr}_{E|F}(ab)$.
@@
Given a conjugacy class $\sigma]$ of an involution $\sigma$ in $\operatorname{Gal}(F)$, we have the ordering $P[\sigma]=\{a\in F\mid\sigma(\sqrt{a})=\sqrt{a}\}$. Note that the condition $\sigma(\sqrt{a})=\sqrt{a}$ does not depend on the choice of the square root $\sqrt{a}$ of $a$. We refer to [33, Thm. I.11.3] for more detail. ∎
@@
Let $F$ be a field and $\operatorname{Gal}(F)$ be its absolute Galois group. From the work of Artin and Schreier cited in the introduction, see [4, Satz 4] in particular, it is known that the group $\operatorname{Gal}(F)$, even though profinite, can only have non-trivial elements of finite order if $\operatorname{char}(F)=0$, and all of these elements necessarily have order $2$, i.e., they are involutions. We refer to [29] for an interesting discussion of torsion in absolute Galois groups from a different angle. The following lemma collects useful group-theoretic statements about involutions in absolute Galois groups.
@@
If $\tau$ is a field automorphism of $\operatorname{Fix}(\sigma)$ that fixes $F$, and $a\in\operatorname{Fix}(\sigma)$ with $\tau(a)\not=a$, we have $a<\tau(a)$, say. Then $a<\tau(a)<\tau^{2}(a)<\tau^{3}(a)<\dots$, contradicting the algebraicity of $a$: the $\tau$–orbit of $a$ must be finite. The argument for $a>\tau(a)$ is similar, and we deduce $\tau(a)=a$ for all $a\in\operatorname{Fix}(\sigma)$. This shows $\operatorname{Aut}(\operatorname{Fix}(\sigma)|F)=\{\operatorname{id}\}$.
If an element $\tau\in\operatorname{Gal}(F)$ lies in the centraliser of $\sigma$, then $\tau$ sends $\operatorname{Fix}(\sigma)$ into itself. From what we have shown before, it follows that $\tau$ is the identity on $\operatorname{Fix}(\sigma)$ and lies in $\operatorname{Gal}(\operatorname{Fix}(\sigma))=\langle\sigma\rangle$. The last statement is simply yet another reformulation of the preceding statement. ∎
Later, we shall need the following stronger statement about involutions in absolute Galois groups.
@@
By Artin–Schreier, we already know that every element of the finite subgroup $G$ must have order $1$ or $2$. It follows from group theory that $G$ is an elementary abelian 22–group, that is $G\cong(\mathbb{Z}/2)^{n}$ for some $n$. We show $n=1$ by showing that $\operatorname{Gal}(F)$ cannot contain a subgroup of the form $\mathbb{Z}/2\times\mathbb{Z}/2$: Given an involution $\sigma\not=\operatorname{id}$, Lemma 3.5 says that its centraliser within $\operatorname{Gal}(F)$ is the subgroup $\langle\sigma\rangle$ generated by $\sigma$. Therefore, there can be no other element in the group $\operatorname{Gal}(F)$ that commutes with $\sigma$. ∎
@@
For any field $F$, the set $\operatorname{Inv}(\operatorname{Gal}(F))$ is a closed subspace of the absolute Galois group $\operatorname{Gal}(F)$. Therefore, it is profinite as well.
@@
Let $F$ be a field. The Artin–Schreier quandle $\mathrm{AS}(F)$ of $F$ is the profinite involutory quandle
@@
The Artin–Schreier quandle $\mathrm{AS}(F)$ of a field $F$, just as its absolute Galois group $\operatorname{Gal}(F)$, depends on the choice of an algebraic closure of $F$. If $\omega\colon\Omega_{1}\to\Omega_{2}$ is an FF–isomorphism of algebraic closures of $F$, then $\sigma\mapsto\omega\sigma\omega^{-1}$ is an isomorphism $\operatorname{Aut}(\Omega_{1}|F)\to\operatorname{Aut}(\Omega_{2}|F)$ and this isomorphism restricts to an isomorphism between the Artin–Schreier quandles computed using $\Omega_{1}$ and $\Omega_{2}$, respectively. As usual, we will suppress the choice of an algebraic closure from the notation.
@@
Before we come to examples, let us note the following algebraic result about a particular property of Artin–Schreier quandles.
@@
The ‘if’ direction is part of the quandle axioms, which are evident by now. For the ‘only if’ direction, note that $\sigma\rhd\tau=\tau$ is equivalent to $\sigma\tau=\tau\sigma$, so that $\tau$ must be in the centraliser of $\sigma$. It cannot be the identity, by definition of $\mathrm{AS}(F)$, so the result follows from Lemma 3.5. ∎
@@
A field $F$ is known to be formally real if and only if its absolute Galois group $\operatorname{Gal}(F)$ contains an involution (see [33, Kor. I.6.2], for instance). ∎
@@
Proposition 3.12 implies that it is not interesting—for our purposes—to consider fields that are not formally real. Therefore, unless otherwise stated, we can assume we are dealing with formally real fields.
@@
We know from the work of Artin and Schreier (see again [4, Satz 4]) that a field $F$ is real closed if and only if its absolute Galois group has order $2$. ∎
@@
The field $\mathbb{R}$ of real numbers is real closed, and the unique element in the involutory quandle $\mathrm{AS}(\mathbb{R})=\{\gamma\}$ is complex conjugation $\gamma\colon z\mapsto\overline{z}$ inside the absolute Galois group $\operatorname{Gal}(\mathbb{R})=\operatorname{Aut}(\mathbb{C}|\mathbb{R})=\{\operatorname{id},\gamma\}$ of order $2$.
@@
A bijection exists between the set $\mathrm{AS}(F)$ and the set of real closed subfields containing $F$ of the algebraic closure of $F$ used to define the absolute Galois group. An involution $\sigma$ corresponds to its fixed field $\operatorname{Fix}(\sigma)$.
@@
When the Artin–Schreier quandle $\mathrm{AS}(F)$ contains more than one element, the absolute Galois group $\operatorname{Gal}(F)$ contains two different involutions, say $\sigma\not=\tau$. Consider the morphism $\mathrm{F}_{2}\{s,t\}\to\operatorname{Gal}(F)$ of groups that sends the generator $s$ to $\sigma$ and $t$ to $\tau$, and suppose there was a non-trivial kernel. Recall that the group $\mathrm{F}_{2}\{s,t\}$ is the infinite dihedral group, which contains an infinite cyclic subgroup generated by the product $st$ of the generators, with inverse $ts$, and all other elements are alternating products $ststs\dots$ and $tstst\dots$ of odd length, which are conjugate to their ‘middle’ factor, either $s$ or $t$. The kernel of the morphism above cannot contain $s$ or $t$ by definition, and it cannot contain any other involution, as these are each conjugate to either $s$ or $t$ inside $\mathrm{F}_{2}\{s,t\}$. It follows that the kernel is a subgroup of the infinite cyclic subgroup generated by $st$, so that the image is a finite (dihedral) subgroup of $\operatorname{Gal}(F)$, and it contains two different involutions $\sigma\not=\tau$. But the finite subgroups of $\operatorname{Gal}(F)$ have order at most $2$ by Lemma 3.6—a contradiction. Hence, the morphism $\mathrm{F}_{2}\{s,t\}\to\operatorname{Gal}(F)$ is injective, and $\operatorname{Gal}(F)$ contains infinitely many involutions because this already holds for the infinite dihedral group $\mathrm{F}_{2}\{s,t\}$, and these lie in $\mathrm{AS}(F)$. ∎